# Md Kowsher School of Engineering and Science

### v1

@@ 3.3 Reservoir Transformer (RT)
Our group reservoir efficiently models observed time-series data in dynamical systems and require minimal training data and computing resources (Gauthier et al., 2021) and is ideal for handling long sequential inputs. Here we mainly describe adopting a reservoir for Transformer Vaswani et al. (2017) as one of the most well-known and valuable DNN architectures. However, our methods are adaptable to other DNN model architectures, for whom we also compared results in our experiments.
Figure a: Non-linear readout for RT.
Figure a: Non-linear readout for RT.
Figure b: Ensemble of multiple reservoirs
Figure 1: Description of deep Reservoir Transformer: In Figure (a), the readout obtained from self-attention, combined with the present input $\vec{u}(t+1)$, serves as the input for training the transformer architecture. In Figure (b), the input $\vec{u}(t)$ is processed individually by $Q$ reservoirs. The nonlinear output of each reservoir is later combined to form the ultimate output, labeled as $\vec{o}(t)$.
Figure 1 shows the architecture of our reservoir-enhanced Transformer. The left figure shows the ensemble of reservoirs, and the right figure shows how to use a non-linear readout reservoir to help Transformer handle arbitrarily long inputs.
@@
Our approach enables comprehensive modeling of entire historical data in decision-making. While the Transformer captures temporal dependencies (Equation 1), limited by a window of size $s$, our method incorporates reservoir pooling to overcome this limitation. This permits the consideration of the entire history without a rise in time complexity. By incorporating Equation 10 (ensemble of all $Q$ reservoir readouts), we arrive at:
$\vec{z}(t+1)=\kappa\vec{u}(t+1)+(1-\kappa)\vec{o}(t)$ (11)
$\bar{\vec{y}}(t+1)=\mathcal{M}(\vec{z}(t+1))$ (12)
$\vec{z}(t+1)$ denotes the Transformer $\mathcal{M}$ input with $\kappa$ as a learnable parameter indicating the weights of current multivariate states and reservoir state to predict $\bar{\vec{y}}(t+1)$ by Equation 12.
In the context of multivariate time series prediction, the target label for forecasting up to $q$ future time steps is denoted as $\vec{u}(t+1:t+1+q;v)$, which is contingent upon the observed series $\vec{u}(1:t)$. In the multivariate case, all available features are taken into account with the conditional distribution $Pr(\vec{u}(t+1:t+1+q)|\vec{u}(1:t);\phi)$.
@@ 3.4 Training
We enhance the non-linear readout layers through self-attention mechanisms within ensemble reservoirs. This involves updating these layers and performing element-wise addition of the readouts, resulting in an output denoted as $\vec{o}(t)$. This output is concatenated with the present input features $\vec{u}(t+1)$, yielding $z(t+1)$, which then undergoes processing via the Transformer encoder $M(.)$. Initially, we sequentially train the model $M(.)$ across all time stamps $1:T$ using the training dataset. Subsequently, we fine-tune our models using the validation set to achieve optimal model performance. The model is subjected to testing with varying hyper-parameters such as reservoir size, leaky rate, spectral radius, number of reservoirs, learning rate, attention size, Transformer block, and dropout rate. The choice of these hyper-parameters may differ depending on the specific dataset employed (details in the appendix). For a comprehensive understanding of our training algorithm, refer to the pseudo-code provided in the appendix, along with a detailed explanation of the notation employed. In this work, we use the Huber loss function Meyer (2021) as our objective (Equation 13):
$\mathcal{L}(\vec{y},\bar{\vec{y}})=\frac{1}{T}\sum_{i=1}^{T}\begin{cases}\frac{1}{2}(\vec{y}_{i}-\bar{\vec{y}_{i}})^{2},&\text{if}\ |\vec{y}_{i}-\bar{\vec{y}}|\leq\delta\\ \delta|\vec{y}_{i}-\bar{\vec{y}_{i}}|-\frac{1}{2}\delta,&\text{otherwise},\end{cases}$ (13)
Here $\delta$ is an adjustable parameter that controls where the function change occurs to keep the function differentiable, $T$ is the total number of samples and $\vec{y},\bar{\vec{y}}$ are ground truth and predicted values respectively. In the same way, for time series
@@ 4 Experiments
We show our experimental results of Reservoir Transformer (RT) compared to baselines, including state-of-the-art methods in time-series prediction, such as: Nlinear (Zeng et al., 2022), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021), Informer (Zhou et al., 2021), Pyraformer (Liu et al., 2021), LogTrans (Li et al., 2019), GRIN (Cini et al., 2021), BRITS (Cao et al., 2018), STMVL (Yi et al., 2016), M-RNN (Yoon et al., 2018), ImputeTS (Moritz & Bartz-Beielstein, 2017), and Transformer Vaswani et al. (2017). To conduct experiments on multivariate time series, we adhered to the established methodology outlined by  (Zeng et al., 2022), which includes data splitting, preprocessing, and setting horizon times.”
4.1 Result Analysis
We thoroughly assess our RT method on different time series regression tasks. These tasks include Electricity, Traffic, Weather, ETTh1, ETTh2, ETTm1, ETTm2, ILI, Exchange Rate (Zhou et al., 2021), Air Quality (Cini et al., 2021), Daily website visitors (DWV), Daily Gold Price (DGC), and Daily Demand Forecasting Orders (DDFO), Bitcoin Historical Dataset (BTC). We also test RT’s performance on two-time series classification tasks, i.e., Absenteeism at work (AW) and Temperature Readings from IOT Devices (details in the Appendix A.1). In our empirical evaluation, presented in Table 3, a comparison of multivariate long-term forecasting errors using the Mean Squared Error (MSE) metric was performed across various methods and datasets. For each dataset, a forecasting horizon was specified with ILI having horizons $T\in\{24,36,48,60\}$ and the rest having $T\in\{96,192,336,720\}$. It is evident that the RT consistently outperforms other methods, achieving the lowest MSE in most scenarios. The Transformer-based methods also show competitive performance, with some of their results being underscored for distinction (univariate time forecasting experiments in the Appendix A.5).
Test Chaotic Behavior:
Figure a: Mackey Glass Series.
Figure b: ECG Signal
Figure a: Mackey Glass Series.
Figure b: ECG Signal
Figure c: Lorenz Attractor
Figure 2: The figures (a), (b), and (c) represent chaotic datasets. The blue line illustrates the multi-horizon (800) predictions generated by our model, while the green line depicts the actual ground truth data.
Time Series $D_{2}$
Length/Dataset MGS ECG Lorenz MGS ECG Lorenz
50 0.173 0.301 0.313 3.165 2.824 2.425
100 0.185 0.314 0.331 3.242 2.913 2.515
500 0.206 0.341 0.359 3.452 3.173 2.721
1000 0.225 0.374 0.447 3.531 3.282 2.841
Table 1: Comparison of Time Series Lengths and Correlation Dimensions ($D_{2}$) for Mackey Glass Series (MGS), Electrocardiogram Signal (ECG), and Lorenz Attractor datasets across varying lengths of 50, 100, 500, and 1000. Higher values of $D_{2}$ indicate increased complexity in the underlying dynamics.
In order to investigate the presence of chaotic behavior and assess the predictive capabilities of our RT models, three distinct chaotic datasets were employed: the Mackey Glass Series (MGS), Electrocardiogram Signal (ECG), and Lorenz Attractor. The experiments involved horizons of 50, 100, 500, and 1000, as outlined in Table 1. Our findings indicate that our RT model successfully captures chaotic dynamics across multiple forecasting horizons, shown in Figure 2.
In addition, to show that the datasets we have used are chaotic, we employ the correlation dimension, denoted as $D_{2}$ serves as a reliable measure of chaotic behavior, with higher values reflecting more pronounced chaos. To evaluate the chaotic behavior of various time series datasets  (Pánis et al., 2020), the $D_{2}$ is defined as
$D_{2}=\underset{\epsilon\to 0}{\lim}\frac{\ln{C(\epsilon)}}{\ln{\epsilon}}.$ (14)
The correlation sum $C(\epsilon)$ for some small scalar $\epsilon$ is defined as in (Grassberger & Procaccia, 1983):
$C(\epsilon)=\underset{N\to\infty}{\lim}\frac{2}{N(N-1)}\sum_{i<j}H(\epsilon-|x_{i}-x_{j}|),$ (15)
where $H$ is the Heaviside step function, $N$ is the number of points, and $|x_{i}-x_{j}|$ is the distance between two points . We compute the correlation dimension for all datasets in our experiments using the implementation provided by (Vasquez-Correa, 2015)
Length/Dataset ETTh1 ETTh2 ETTm1 ETTm2 Exchange ILI BTC DGP DWV DDFO
100 0.790 1.190 0.560 0.299 0.995 0.995 0.961 1.070 1.343 0.540
500 1.340 0.973 0.931 0.535 1.099 1.249 0.853 1.626 1.823 0.540
1000 1.266 0.866 1.061 0.667 1.359 1.299 0.853 1.398 1.838 0.540
Table 2: This study investigates the correlation dimension $D_{2}$ across multiple datasets using diverse state numbers (100, 1000, and 10000). The correlation dimension $D_{2}$ serves as an indicator of chaotic behavior, with higher values denoting a more pronounced presence of chaos (Panis, 2020). Our findings indicate that the correlation dimension $D_{2}$ was evaluated for all datasets under varying state numbers.
Table  1 and  2 display the outcomes of a correlation dimension ($D_{2}$) analysis across various datasets. The correlation dimension gauges chaotic tendencies, with higher values suggesting heightened chaos. This assessment involves different state numbers: 50, 100, 500, and 1000 and it exhibits $D_{2}$ values for each dataset and state number. The findings reveal that, for all datasets, the correlation dimension increases with higher state numbers, indicating intensified chaotic behavior. Since Mackey Glass Series (MGS), Electrocardiogram Signal (ECG), and Lorenz Attractor datasets are highly chaotic datasets, they showed the higher $D_{2}$.
Model
Horizons
Linear
NLinear
@@ 4.1 Result Analysis
Autoformer
Informer
Pyraformer
LogTrans
Repeat

### v2

@@ 3.7 Group Reservoir Transformer (RT)
Our group reservoir efficiently models observed time-series data in dynamical systems and require minimal training data and computing resources (Gauthier et al., 2021) and is ideal for handling long sequential inputs. Here we mainly describe adopting a reservoir for Transformer as one of the most well-known DNN architectures. However, our methods are adaptable to other DNN model architectures, for whom we also compared results in our experiments.
Figure 1: Left: Group Reservoir Transformer; Right: Our RT prediction is closer to the ground truth than the Transformer in Lorenz Extractor.
Figure 1 (Left) shows the architecture of our reservoir-enhanced Transformer. We enhance the non-linear readout layers through self-attention mechanisms within ensemble reservoirs. This involves updating these layers and performing element-wise addition of the readouts, resulting in an output denoted as $\mathbf{z}_{t}$. This output is concatenated with the present input embedding $\mathbf{h}_{t}$, yielding $\mathbf{f}_{t}$, which then undergoes processing via the Transformer encoder $\mathbb{M}(.)$.
@@
Our approach enables comprehensive modeling of entire historical data in decision-making. While the Transformer captures temporal dependencies, limited by a window of size $w$, our method incorporates reservoir pooling to overcome this limitation. This permits the consideration of the entire history without a rise in time complexity. By incorporating Equation 7 (ensemble of all $L$ reservoir readouts) and then following Equation 10, we have:
$\mathbf{f}_{t}=(1-\kappa)\mathbf{z}_{t}\uplus\kappa\mathbf{h}_{t}$ (11)
$\mathbf{\hat{u}}_{t+1:t+\tau}=\mathbb{M}(\mathbf{f}_{t})$ (12)
[…]
@@ 3.8 Training
In this work, we use the Huber loss function (Meyer, 2021) as our objective (Equation 13):
$\mathcal{L}(\mathbf{u}_{t+1:t+\tau},\mathbf{\hat{u}}_{t+1:t+\tau})=\frac{1}{\tau}\sum_{i=t+1}^{t+\tau}\begin{cases}\frac{1}{2}(\mathbf{u}_{i}-\mathbf{\hat{u}}_{i})^{2},&\text{if }|\mathbf{u}_{i}-\mathbf{\hat{u}}_{i}|\leq\delta\\ \delta(|\mathbf{u}_{i}-\mathbf{\hat{u}}_{i}|-\frac{1}{2}\delta),&\text{otherwise}\end{cases}$ (13)
Here $\delta$ is an adjustable parameter that controls where the function change occurs to keep the function differentiable. For a comprehensive understanding of our training algorithm, refer to the pseudo-code provided in the Appendix B, along with a detailed explanation of the notation employed. Initially, we sequentially train the model $\mathbb{M}(.)$ across all $T$ time stamps using all the training dataset. Subsequently, we fine-tune our model using the validation set to achieve optimal model performance. The model is subjected to testing with varying hyper-parameters such as reservoir size, leaky rate, spectral radius, number of reservoirs, learning rate, attention size, Transformer block, and dropout rate. The choice of these hyper-parameters may differ depending on the specific dataset with details in Appendix A.
@@ 4 Experiments
Baselines: We use well-known time-series forecasting models for our baseline comparisons, including Transformer based systems such as DLinear (Zeng et al., 2023), PatchTST (Nie et al., 2022), TimesNet (Wu et al., 2022a), FEDformer (Zhou et al., 2022), Stationary (Liu et al., 2022), ESTformer (Woo et al., 2022) , LightTS (Campos et al., 2023), and Autoformer (Chen et al., 2021), and pre-trained LLMs-based systems such as TimeLLM (Jin et al., 2024) and GPT2TS (Zhou et al., 2024). The experimental setup follows  (Zeng et al., 2023) for all baseline models.
Datasets: We evaluate RT on several datasets commonly used for benchmarking long-term multivariate forecasting models, including four ETT datasets (ETTh1, ETTh2, ETTm1, and ETTm2) (Zhou et al., 2021), Weather, Electricity (ECL), Traffic, and ILI from Zeng et al. (2022). To test the chaotic behavior on these datasets, we employ the correlation dimension, denoted as $D_{2}$ to serve as a reliable measure of chaotic behavior, with higher values reflecting more pronounced chaos (Pánis et al., 2020). We compute correlation dimension $D_{2}$ across multiple datasets using diverse state numbers ($100$, $500$, and $1000$). Table 1 shows the $D_{2}$ of each dataset.
Length ETTh1 ETTh2 ETTm1 ETTm2 DDFO Exchange ILI BTC DGP DWV MGS ECG LA
100 0.790 1.190 0.560 0.299 0.540 0.995 0.995 0.961 1.070 1.343 3.242 2.913 2.515
500 1.340 0.973 0.931 0.535 0.540 1.099 1.249 0.853 1.626 1.823 3.452 3.173 2.721
1000 1.266 0.866 1.061 0.667 0.540 1.359 1.299 0.853 1.398 1.838 3.531 3.282 2.841
low chao medium chao high chao
Table 1: $D_{2}$ chaotic behavior measure: MGS, ECG, LA are high chaos; ETTh are low chaos.
Results: The time-series forecasting prediction error rate results are presented in Table 2. The input time series length $k$ is set to 300, and we test four different prediction horizons $h\in\{96,192,336,720\}$. The evaluation metrics are mean square error (MSE) and mean absolute error (MAE). Our RT consistently outperforms all baselines in most cases, achieving significant error reductions. Specifically, compared to DLinear, RT reduces MSE by an average of 4.2%. Similarly, the reductions in MSE for PatchTST and TimeNet are 1.4% and 2.4%, respectively. Additionally, RT outperforms pre-trained LLM-based time series models in Table 3. Table 4 shows that RT model outperforms baseline Transformer on various datasets: Air Quality (AQ), Daily Website Visitors (DWV), Daily Gold Price (DGP), Daily Demand Forecasting Orders (DDF), Bitcoin Historical Dataset (BTC), Absenteeism at Work (AW), and Temperature Readings (TR). All results are reported in MSE and MAE values, with only AW and TR reporting Accuracy and F-Score metrics. Comparisons with more baselines in other evaluation criteria are detailed in Appendix F.
Methods
RT
@@ Methods
Stationary
ETSformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.362 0.392 0.376 0.397 0.375 0.399 0.370 0.399 0.384 0.402 0.376 0.419 0.449 0.479
192 0.396 0.412 0.405 0.416 0.413 0.421 0.436 0.429 0.420 0.448 0.534 0.504 0.538 0.504
336 0.427 0.422 0.439 0.443 0.422 0.436 0.491 0.469 0.459 0.465 0.588 0.535 0.574 0.521
720 0.441 0.455 0.472 0.490 0.447 0.466 0.521 0.500 0.506 0.507 0.643 0.616 0.562 0.535
avg 0.406 0.418 0.422 0.437 0.413 0.430 0.458 0.450 0.440 0.460 0.570 0.537 0.542 0.510
ETTh2 96 0.262 0.320 0.289 0.353 0.274 0.336 0.340 0.374 0.358 0.397 0.476 0.458 0.340 0.391
192 0.331 0.373 0.383 0.418 0.339 0.379 0.402 0.414 0.429 0.439 0.512 0.493 0.430 0.439
336 0.358 0.403 0.448 0.465 0.329 0.380 0.452 0.452 0.496 0.487 0.552 0.551 0.485 0.479
720 0.369 0.411 0.605 0.551 0.379 0.422 0.462 0.468 0.463 0.474 0.562 0.560 0.500 0.497
avg 0.330 0.376 0.431 0.446 0.330 0.379 0.414 0.427 0.439 0.452 0.602 0.543 0.450 0.459
ETTm1 96 0.272 0.336 0.299 0.343 0.290 0.342 0.338 0.375 0.379 0.419 0.386 0.398 0.375 0.398
192 0.310 0.357 0.335 0.365 0.332 0.369 0.374 0.387 0.426 0.441 0.459 0.444 0.408 0.410
336 0.351 0.382 0.369 0.386 0.366 0.392 0.410 0.411 0.445 0.459 0.495 0.464 0.435 0.428
720 0.380 0.409 0.425 0.421 0.416 0.420 0.478 0.450 0.543 0.490 0.585 0.516 0.499 0.462
avg 0.328 0.371 0.357 0.378 0.351 0.380 0.400 0.406 0.448 0.452 0.481 0.456 0.429 0.425
ETTm2 96 0.163 0.254 0.167 0.269 0.165 0.255 0.187 0.267 0.203 0.287 0.192 0.274 0.189 0.280
192 0.218 0.290 0.224 0.303 0.220 0.292 0.249 0.309 0.269 0.328 0.280 0.339 0.253 0.319
336 0.271 0.328 0.281 0.342 0.274 0.329 0.321 0.351 0.325 0.366 0.334 0.361 0.314 0.357
720 0.356 0.382 0.397 0.421 0.362 0.385 0.408 0.403 0.421 0.415 0.417 0.413 0.414 0.413
avg 0.252 0.627 0.267 0.333 0.255 0.315 0.291 0.333 0.305 0.349 0.371 0.306 0.347 0.293
Weather 96 0.146 0.196 0.176 0.237 0.149 0.198 0.172 0.220 0.217 0.296 0.173 0.223 0.197 0.281
192 0.193 0.236 0.220 0.282 0.194 0.241 0.219 0.261 0.276 0.336 0.245 0.285 0.237 0.312
336 0.239 0.240 0.265 0.319 0.245 0.282 0.280 0.306 0.339 0.380 0.321 0.338 0.298 0.353
720 0.301 0.316 0.333 0.362 0.314 0.334 0.365 0.359 0.403 0.428 0.414 0.410 0.352 0.288
avg 0.219 0.247 0.248 0.300 0.225 0.264 0.259 0.287 0.309 0.360 0.288 0.314 0.271 0.334
Electricity 96 0.130 0.219 0.140 0.237 0.129 0.222 0.168 0.272 0.193 0.308 0.169 0.273 0.187 0.304
192 0.154 0.243 0.153 0.249 0.157 0.240 0.184 0.289 0.201 0.315 0.182 0.286 0.199 0.315
336 0.158 0.245 0.169 0.267 0.163 0.259 0.198 0.300 0.214 0.329 0.200 0.304 0.212 0.329
720 0.190 0.294 0.203 0.301 0.197 0.290 0.220 0.320 0.246 0.355 0.222 0.321 0.233 0.345
avg 0.158 0.251 0.166 0.263 0.161 0.252 0.192 0.295 0.214 0.327 0.193 0.296 0.208 0.323
Traffic 96 0.361 0.248 0.410 0.282 0.360 0.249 0.593 0.321 0.587 0.366 0.612 0.338 0.607 0.392
192 0.376 0.250 0.423 0.287 0.379 0.256 0.617 0.336 0.604 0.373 0.613 0.340 0.621 0.399
336 0.382 0.268 0.436 0.296 0.392 0.264 0.629 0.336 0.621 0.383 0.618 0.328 0.622 0.396
720 0.426 0.281 0.466 0.315 0.432 0.286 0.640 0.350 0.626 0.382 0.653 0.355 0.632 0.396
avg 0.386 0.261 0.433 0.295 0.390 0.263 0.620 0.336 0.610 0.376 0.628 0.379 0.624 0.340
ILI 24 1.301 0.739 2.215 1.081 1.319 0.754 2.317 0.934 3.228 1.260 2.294 0.945 2.527 1.020
36 1.409 0.824 1.963 0.963 1.430 0.834 1.972 0.920 2.679 1.080 1.825 0.848 2.615 1.007
48 1.520 0.807 2.130 1.024 1.553 0.815 2.238 0.940 2.622 1.078 2.010 0.900 2.359 0.972
60 1.553 0.852 2.368 1.096 1.520 0.788 2.027 0.928 2.857 1.157 2.178 0.963 2.487 1.016
avg 1.439 0.805 2.169 1.041 1.443 0.797 2.139 0.931 2.847 1.144 2.077 0.914 2.497 1.004
Table 2: Multivariate long-term time-series forecasting: lower values indicate better performance. Forecasting horizons are $h\in\{24,36,48,60\}$ for ILI and $h\in\{96,192,336,720\}$ for the rest.
Methods